# Consensus formation in quality-sensitive interdependent agent systems

### v2

@@ A.3 Imperfect quality assessment: The case of non-null parameter $\mu$
As we explained in the main text, the parameter $\mu$ in the orginal LES model governs the bees’ quality assessment behavior. When $\mu\rightarrow 0$, bees independently assess the quality of sites, and thus, the duration of the advertisement period strongly correlates with the actual site quality. Conversely, when $\mu\rightarrow 1$, bees do not individually assess site qualities but rather imitate their peers without regard to the options’ qualities. Consequently, the advertisement duration is influenced by a generic parameter $K$ that is the same for all sites, as indicated in Eq. (3). Although we have primarily considered $\mu=0$ in the text to explore the interplay between other parameters such as interdependence, discovery probabilities, and qualities, here we provide a brief overview of how the results change when nonzero values of the independent quality assessment parameter, $\mu$, are considered. For simplicity and in accordance with the prescription by List et al. in the original formulation of the model [21], we set $K$ equal to the maximum quality among all sites, i.e., $K=q_{k}$.
@@ A.5 Lattice simulations in $d=2$
We additionally implement a spatially distributed version of the model described in Sec II. In the spatially distributed version of the model, each scout bee occupies a specific location on a square lattice with periodic boundary conditions. The lattice has a lateral length of $L$, resulting in a total of $N=L\times L$ nodes. Each scout bee interacts only with its nearest neighbors on the lattice.
The state $s_{i}$ of each bee $i$ now corresponds to a specific location ${\bf r}=(x,y)$ on the lattice, denoted as $s_{i}(t)\equiv s({\bf r},t)\equiv s(x,y;t)$. The fractions $f_{\alpha,t}$ of bees promoting site $\alpha$ at time $t$, which enter into the equation defining the system dynamics (Eq. 2), are computed as follows:
@@
where $\delta$ is the Kronecker’s delta, the sum runs over the four nearest neighbors of ${\bf r}$ on the square lattice, and the index $\alpha=1,\ldots,k$ indicates the possible committed states. In addition to the spatial dependence of $f_{\alpha}$, the state update rules at each site, as described by Eq. (II), remain the same. Each time step involves updating (or attempting to update) the states of every bee or node in the lattice. We implement this process using a massively parallel programming approach and ensure that the evolution of interdependent nodes on each time step is serialized by using a checkerboard lattice decomposition.
The spatially distributed model, similar to [31], is not designed to replicate realistic spatial behavior of bees but rather to analyze the dynamic competition of states within a spatio-temporal framework. As discussed in the main text, the general stationary results of this spatial model align perfectly with mean field (i.e., fully-connected) simulations. Analyses such as the time to reach the stationary state, as depicted in Fig. 4F or Fig. 6E, also yield qualitatively similar results, exhibiting the same trends discussed in the main text - see Fig. A3.

### v3

@@ A.3 Imperfect quality assessment: The case of non-null parameter $\mu$
As we explained in the main text, the parameter $\mu$ in the orginal LES model governs the bees’ quality assessment behavior. When $\mu\rightarrow 0$, bees independently assess the quality of sites, and thus, the duration of the advertisement period strongly correlates with the actual site quality. Conversely, when $\mu\rightarrow 1$, bees do not individually assess site qualities but rather imitate their peers without regard to the options’ qualities. Consequently, the advertisement duration is influenced by a generic parameter $K$ that is the same for all sites, as indicated in Eq. (3). Although we have primarily considered $\mu=0$ in the text to explore the interplay between other parameters such as interdependence, discovery probabilities, and qualities, here we provide a brief overview of how the results change when nonzero values of the independent quality assessment parameter, $\mu$, are considered. For simplicity and in accordance with the prescription by List et al. in the original formulation of the model List et al. (2008), we set $K$ equal to the maximum quality among all sites, i.e., $K=q_{k}$.
@@ A.5 Lattice simulations in $d=2$
We additionally implement a spatially distributed version of the model described in Sec II, as a cellular automaton. In this spatially distributed version of the model, each scout bee occupies a specific location on a square lattice with periodic boundary conditions. The lattice has a lateral length of $L$, resulting in a total of $N=L\times L$ nodes. Each scout bee interacts only with its nearest neighbors on the lattice.
The system is defined by a scalar field $\{s\}$ where $s_{i}$ ($i=1,\ldots,N$) stands for the state of each bee $i$ and corresponds to a specific location ${\bf r}=(x,y)$ on the lattice, denoted as $s_{i}(t)\equiv s({\bf r},t)\equiv s(x,y;t)$. Starting from an initial configuration $\{s\}(t=0)$, the system evolves according to the update rules defined in Eqs. (II) and  (2). The fractions $f_{\alpha,t}$ of bees promoting site $\alpha$ at time $t$ entering in Eq. (2), are now computed in the square lattice as:
@@
where $\delta$ is the Kronecker’s delta, the sum runs over the four nearest neighbors of ${\bf r}$ on the square lattice, and the index $\alpha=1,\ldots,k$ indicates the possible committed states. The state update rules at each site, as described by Eq. (II), remain the same. A time step consist in updating every local state $s_{i}$ probabilistically (Eq. (II)). As a remark, we computationally accelerate this update process by using a massively parallel approach and GPUs as a hardware. The technical approach to ensuring the correctness of the parallel implementation involves using a checkerboard lattice decomposition. This method alternates between updating all black cells in parallel and all white cells in parallel, preventing the overwriting of state values before they are used to compute transition probabilities. This parallelization technique ensures consistency with a serial implementation, where each cell is updated one at a time, maintaining equivalence at the global step balance level. It has been proven effective in various simulations, such as Monte Carlo Ferrero et al. (2012) and Ginzburg-Landau Kolton et al. (2023) simulations. Additionally, observables such as population fractions, the order parameter, and fluctuations are calculated using parallel reduction algorithms.
The spatially distributed model, similar to Galla (2010), is not designed to replicate realistic spatial behavior of bees but rather to analyze the dynamic competition of states within a spatio-temporal framework. As discussed in the main text, the general stationary results of this spatial model align perfectly with mean field (i.e., fully-connected) simulations. Analyses such as the time to reach the stationary state, as depicted in Fig. 4F or Fig. 6E, also yield qualitatively similar results, exhibiting the same trends discussed in the main text - see Fig. A3.